OpenStudy (anonymous):

Show that every normal line to th sphere x^2+y^2+z^2 = r^2 where r is the radius passes through the center of the sphere.

OpenStudy (experimentx):

take gradient of the surface

OpenStudy (anonymous):

ok. (2x,2y,2z) is it?

OpenStudy (experimentx):

yeah ... the normal is x - x1 ..... ----- = -------- 2x_1 .....

OpenStudy (experimentx):

put x_1, y_1, z_1 in place of x, y, z ... assume it to be point on surface of sphere.

OpenStudy (anonymous):

i need guidance

OpenStudy (experimentx):

or in parametric form \[ r(t) = (x_1.y_1,z_1)+t(2x_1.2y_1,2z_1) \] for t=-1/2 ... this always pass through (0,0,0)

OpenStudy (experimentx):

or just put the line in this form